Fermi National Laboratory


Accelerator Update for July 18, 2005 - July 20, 2005

Monday July 18
The day shift began with Operations monitoring Tevatron (TeV) store 4279 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following colliding, neutrino, and fixed target experiments taking beam: the Collider Detector at Fermilab (CDF), D-Zero (D0), MiniBooNE, NuMI, and the Meson MCenter experiment.

At 8:43 AM, Operations reported that a Main Injector (MI) RF station (MIRF5) tripped off for the first of four times in thirty-six minutes. An expert investigated after the last trip.

At 8:47 AM, Operations reported that MiniBooNE was down due to a quadrupole power supply (E:Q874) tripping off. An expert investigated and switched the power supply to a different breaker. E:Q872 was switched to the weak breaker.

Operations terminated store 4279 at 11:02 AM and then put the TeV into shot setup.

At 1:39 PM, Linac experts took an RF station (LRF2) offline to repair a LCW leak. This halted all downstream beam.

Operations established combination store 4280 at 2:31 PM with an initial luminosity of 105.81E30.

Linac experts completed their repair of LRF2 at 2:48 PM.

Operations established beam to MiniBooNE, NuMI, and MCenter.

Pbar resumed stacking at 3:09 PM.

Pbar experts began a Debuncher stacking study at 3:56 PM.

At 4:27 PM, Operations reported that NuMI had tripped off for the first of six times in three hours due to beam in the kicker gap.

ECool experts inserted their collimators for a study 9 PM. There were no stored antiprotons in the Recycler.

At 9:50 PM, the D0 experimenters reported a problem with beam halo. Operators investigated and discovered a sector D0 quadrupole's position (roll) had changed (D0Q4).

Tuesday July 19
The midnight shift began with Operations monitoring store 4280, with Pbar conducting stacking studies, and with MiniBooNE, NuMI, and MCenter.

At 12:13 AM, a MiniBooNE experimenter asked to have their beam turned off while they investigated a problem with their DAQ.

Pbar studies ended and stacking resumed at 1:40 AM.

NuMI reported at 2:11 AM that they had solved their DAQ problems. They resumed taking beam.

At 7:01 AM, the TeV suffered an abort due to a RF supply (TRF5) tripping off. The store was lost. There was no quench. RF experts investigated and began replacing a modulator. They also decided to replace a weak power amplifier for TRF1.

Operations prepared the CDF and D0 collision halls for access 8:22 AM.

Operations put the Recycler into shot setup at 9:09 AM to stash antiprotons. The stash ended at 10:14 AM. Two shots were lost due to a problem with a Debuncher Lambertson (D:ELAM). An expert investigated and discovered bad power lead.

Pbar resumed stacking at 10:36 AM.

TeV RF expert reported at 10:41 AM that they had finished their RF work.

Operations established beam to MiniBooNE at 10:54 AM.

CDF came out of access at 11:44 PM.

Operations put the TeV through a dry squeeze at 12:14 PM.

At 6:07 PM, MiniBooNE tripped off due to E:Q874 tripping off. An expert suggested that the power leads get switched again. Operations called in a power supply expert.

At 6:43 PM, Operations reported that MI beam was having problems due to low-level RF and dampers. Experts tried various reboots with only partial success. An operator called in an expert who restored the damper code.

TeV experts completed their studies at 7:29 PM.

Operations put the TeV into shot setup at 8:26 PM.

The power lead swap between E:Q874 and E:Q872 was completed at 8:20 PM.

At 9:53 PM, Operations reported problems with MI coalescing. After talking with experts, operators changed back to the old damper control board for the shot. This unfortunately, turned off NuMI.

Operations established combination store 4285 at 11:46 PM with an initial luminosity of 68.76E30. Pbar resumed stacking.

Operations switched back to the new MI damper board at 11:52 PM.

MCenter resumed taking beam at 11:56 PM.

Wednesday July 20
The midnight shift began with Operations monitoring store 4285, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began stashing antiprotons at 5:30 AM.

The Recycler stash ended and Pbar resumed stacking at 5:57 AM.

At 7:59 AM, CDF experimenters reported that they were having problems with their solenoid and had to ramp it down.

The Plan for Wednesday
The plans for today are to monitor store 4285 and conduct some machine studies. The rest of the time will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/20/2005   email Fermilab